# Looking through the Kerr disk

### v1

@@ V.1.3 $\theta$-integral
With the solution for the polar angle, we proceed similarly to the previous subsection and state the final formula derived in [gralla_null_2020] (the details can be found in Sec. III B there). The polar angle along the geodesic is given by
@@ V.1.4 $\psi$-integral
We can follow [gralla_null_2020] and rewrite (52) as
@@
$\mathcal{I}_{reg,\varphi}=\nu_{r}\,\frac{a\,(\mathrm{ln}(r-r_{+})-\mathrm{ln}(r-r_{-}))}{r_{+}-r_{-}}\,.$ (56)
Furthermore, as shown in [gralla_null_2020], (54) equates to
@@
and where $\gamma_{\pm}$ and $\mathcal{G}_{\theta}(\theta)$ are given by (46) and (47), respectively. However, we could not confirm the antiderivative of (38) presented in [gralla_null_2020] when cross-checking it with numerical integration. Thus, we calculated $\mathcal{I}_{\pm}$ anew with the help of Mathematica [wolfram_research_inc_mathematica_nodate]. The result can be found in (116)-(118) in Appendix E. The azimuthal angle of the source $\psi_{s}$ can be computed using (51) together with (56), (57) and (116). An example azimuthal trajectory can be found in Fig. 6.
@@ V.1.5 $u$-integral
As before, we can rewrite $I_{t}$ as in [gralla_null_2020] to get
@@
$\mathcal{I}_{reg,t}(r)=$ $\nu_{r}\Big(r+\frac{a^{2}+r_{-}^{2}}{r_{-}-r_{+}}\,\mathrm{ln}(r-r_{-})$
@@
Next, we can use the solution of (63) as presented in [gralla_null_2020], for which they found
@@
Lastly, we need the antiderivative of (39) for $\ell=0,1,2$. For $\ell=0$, the solution is (42). For $\ell=1,2$, respective formulas are presented in [gralla_null_2020]. However, we could not confirm these results, and consequently, we calculated $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ anew. The results are presented in (117) and (118) in Appendix E. An example trajectory using (60) can be found in Fig. 7.
@@ V.2 Numerical solution
To solve this system numerically, we impose six initial conditions (two for each second-order equation and one for each first-order equation). The four initial values are the positional coordinates $(u_{o},\tilde{r}_{o},\theta_{o},\psi_{o})$ of the observer. We set the Mino time $\tau=0$ at the observer, thus we fix $\tilde{r}(0)=-\pi/2$, $\theta(0)=\theta_{o}$, $\psi(0)=\psi_{o}$, $u(0)=u_{o}$. For each second-order equation, we additionally need to state the rate of change of the coordinate at the observer, given by $\tilde{r}^{\prime}(0)=\nu_{r}$ and $\theta^{\prime}(0)=\beta$. We perform the numerical integration using Mathematica’s implementation of the explicit Bogacki-Shampine adaptive Runge-Kutta method [wolfram_research_inc_mathematica_nodate]. Because we start the integration at the observer at $\tilde{r}_{o}=-\pi/2$, we will stop it once we reach the source at $\tilde{r}_{s}=\pi/2$. A comparison of the numerical and analytical solutions is presented in Appendix F.
@@ VI Visualizations
Using the presented solutions to the geodesic equations, we can inspect various vortical null geodesics in more detail. In the following, we consider a black hole with a rotation parameter $a/m=0.99$ and integrate from the position of the observer $(u_{o}=0,\,r_{o}=-\infty,\,\theta_{o}=\pi/4,\,\psi_{o}=0)$ to a source located at $r_{s}=+\infty$. These values are chosen so that the inner throat remains significantly large while maximizing the asymmetry due to the rotation of the central object. Note that by inverting the parameter $\tau$ along the geodesic, we obtain a trajectory that originates at the source and terminates at the observer.
In this section, for presentation, we use Kerr-Schild coordinates, which have an underlying flat metric, and we treat the coordinates $(x,y,z)$ as pseudo-Cartesian coordinates. Furthermore, going from one asymptotically flat region to the other (i.e., from $r>0$ to $r<0$ or vice versa) requires traversing the disk enclosed by the ring singularity [chrusciel_structure_2020]. In Kerr-Schild coordinates, the vortical null geodesics we consider can cross over from positive to negative values of $z$ only by traversing the disk in the $x$-$y$-plane, thereby changing the sign of $r$. Thus, in the following plots, the region with positive/negative $z$ corresponds to the region with positive/negative radius $r$. These two regions correspond to the northern hemisphere in their respective asymptotically flat ends of the spacetime (defined by the polar angle $\theta\in[0,\pi/2]$). Consequently, the horizons are only drawn at positive $z$ while the causality-violating region is only present at negative $z$.
@@ VI.1.2 Geodesics close to the inner throat boundary
Here, we address null geodesics close to the inner throat boundary (approximately $10^{-10}$ away). A geodesic close to the left vertex is shown in Fig. 9. Its main feature is that it remains close to the ring singularity in the region with negative $r$ for a significant portion of its journey. This period features several revolutions around the axis of symmetry and multiple polar oscillations between $\theta_{+}$ and $\theta_{-}$. However, it never crosses the causality-violating region.
For the right vertex of the inner throat, a corresponding trajectory in Kerr-Schild coordinates is presented in Fig. 10. The $\psi$-coordinate of this geodesic is not monotonically increasing or decreasing. In the exterior of a Kerr black hole, the order of null geodesics can be defined as the number of times it crosses the equatorial plane from the emitter to the observer [cunningham_optical_1972]. Equivalently, one can consider the number of turning points in the polar angle along the trajectory [james_gravitational_2015]. Only the latter of these two definitions is meaningful for vortical null geodesics, which can never cross the equatorial plane. With this, the geodesic in Fig. 9 is of order four, while the one in Fig. 10 is of order three.
Furthermore, the geodesic close to the right vertex comes close to the ring singularity (up to $10^{-4}m$) but can never hit it [chrusciel_structure_2020]. However, it traverses the causality-violating region. To see this, we inspect the value of the metric component $g_{\psi\psi}$ along the geodesic. Fig. 11 shows the resulting plot. Since there are regions with $g_{\psi\psi}<0$, this geodesic passes through the causality-violating region multiple times along its path.
@@ VI.2.1 Field of view
With our solutions to the equations of motion for vortical null geodesics, we can furthermore answer the question of what portion of the sky at $r_{s}=+\infty$ is visible for an observer at $(u_{o}=0,\,r_{o}=-\infty,\,\theta_{o}=\pi/4,\,\psi_{o}=0)$. To this end, we discretize the inner throat, resulting in almost $1.5\cdot 10^{6}$ pairs of impact parameters $(\alpha,\,\beta)$ with the closest distance to the inner throat boundary of $3\times 10^{-4}$. We calculate $\theta_{s}$ and $\psi_{s}$ for each of these parameters using our analytic solutions. To better track these pairs, we color-code them based on their distance to the center of the inner throat – warm colors are in the middle, while cool colors correspond to the region close to the inner throat boundary (see Fig. 12(a)).
@@
A further feature of the resulting curves in the polar plot is that the curves with a constant distance to the inner throat boundary develop caustic-like structures before they fold in on themselves at specific distances. Due to the overlap of curves, this is not visible in Figs. 12(b) and 12(c). However, it can be observed in a short animation available at [noauthor_httpswwwquantagonatmasters-thesis_nodate_www]. We decrease the distance to the inner throat from frame to frame, thus approaching the boundary until the analytic and numerical solutions cease to coincide up to an acceptable margin (cf. Appendix F). It is visible that the field of view shifts in the anticlockwise direction in the polar plot as the corresponding curve in the inner throat approaches the boundary, thus covering large portions of the sky at $r_{s}=+\infty$. The fact that the curve in the polar plot folds in on itself results in the field of view covering the same part of the sky more than once. Consequently, the observer perceives multiple images of the same radiating object at $r_{s}=+\infty$ (as in the exterior case of a Kerr black hole [james_gravitational_2015]).
@@ VI.2.2 Observing the sky
A similar question concerns the appearance of a fixed light source at $r_{s}=+\infty$ (e.g., the celestial sphere around the black hole) as seen by an observer at $(u_{o}=0,\,r_{o}=-\infty,\,\theta_{o}=\pi/4,\,\psi_{o}=0)$. For an initial study of the distortion effects (rotation, reflection, stretching, or compression of the image), we consider two sample light sources illuminating the sky at $r_{s}=+\infty$. We argue in Appendix G that we do not need to consider any gravitational redshift when the photons are emitted at $r_{s}=+\infty$ and observed at $r_{o}=$-$\infty$.
@@ VII Summary
An analysis of the angular potential showed that there are, at most, two distinct geodesics with a constant polar angle inside the inner throat. One is always visible and corresponds to the principal null direction on which the Eddington-Finkelstein-like coordinates are based. As such, it also has constant $\psi$ and $u$. The other one is only inside the inner throat for specific values of the rotation $a/m$ and the observer’s polar angle $\theta_{o}$. Moreover, vortical null geodesics can never cross the equatorial plane $\theta=\pi/2$. Consequently, an observer at $r_{o}=$-$\infty$ will only see at most half of the celestial sphere at $r_{s}=+\infty$. In addition, there is a further forbidden region around the equatorial plane from which such an observer receives no photons, see Fig. 3 and [noauthor_httpswwwquantagonatmasters-thesis_nodate].
To visualize individual trajectories and simulate the observer’s view, we solved the geodesic equations of motion for vortical null geodesics inside the inner throat in Eddington-Finkelstein-like coordinates analytically and numerically, with the numerical integration providing an independent validation of the analytical solution. This way, we could correct some formulas in [gralla_null_2020].
@@ Acknowledgements.
We thank Piotr Chruściel for valuable discussions and comments. This work is based on the Master’s thesis of the second author [noauthor_httpswwwquantagonatmasters-thesis_nodate]. The first author acknowledges the support of the Austrian Science Fund (FWF) through Project P 36455 and the START-Project Y 963.
@@ Appendix A $\psi$ and $u$ for Principal Null Geodesics
As a geodesic with $(\alpha=+a\sin\theta_{o},\,\beta=0)$ corresponds to the principal null congruence on which the transformation to Eddington-Finkelstein-like coordinates is based on (cf. [poisson_relativists_2004]), we claim that along this geodesic also $\psi$ and $u$ remain constant, which we prove below. Let us begin by evaluating the constants of motion (23) at these impact parameters, leading to
@@ Appendix B Definitions for the antiderivatives
To calculate the antiderivatives in Sec. V.1, we define additional quantities similar to [gralla_null_2020].
In the case of no radial turning point, all roots of the radial potential $R(r)$ are complex with $r_{1}=r_{2}^{*}$ and $r_{3}=r_{4}^{*}$ (where the star denotes complex conjugation). These roots are given by [gralla_null_2020]
@@
$h$ $=\mathrm{sign}(\cos\theta_{s})\,.$ (100)
For the calculation of the radial solution, we additionally need
@@ Appendix C Elliptic Integrals
For completeness, this section provides a basic introduction to elliptic functions following [abramowitz_handbook_1972] and [gradshtein_table_2000]. An integral of the form $\int R(x,y)\,\mathrm{d}x$ is called an elliptic integral if $R(x,y)$ is a rational function and $y=\sqrt{P(x)}$ where $P(x)$ is a polynomial either cubic or quartic in $x$. Generally, these integrals cannot be solved using elementary functions. All elliptic integrals can be reduced to integrals of elementary functions and three canonical elliptic integrals:
@@ Appendix E Antiderivatives of $I_{\pm}$, $I_{1}$, $I_{2}$
Here, we present the antiderivatives $I_{\pm}$, $I_{1}$, and $I_{2}$, defined in (38) and (39). We calculated these anew because we could not confirm the formulas presented in [gralla_null_2020] when comparing them with numerical methods.
@@ Appendix F Comparing analytics and numerics
Far inside the inner throat (up to a distance of $10^{-3}$ from the inner throat boundary), the difference between the analytical and the numerical solutions for $\theta_{s}$ and $u_{s}$ is less than $10^{-9}$, while for $\tau_{s}$ and $\psi_{s}$ it is less than $10^{-8}$. We interpret this as a complete agreement of the two solutions. Fig. 14 shows the difference of the numerical and analytic solution along a vortical null geodesic with impact parameters $(\alpha=-0.15,\,\beta=0.1)$, which corresponds to the example presented in the figures in Sec. V.1. As expected, the error grows with decreasing distance to the throat boundary. At a distance of $10^{-8}$ from the boundary the maximal error for $\theta_{s}$ and $u_{s}$ is around $10^{-5}$, and $10^{-4}$ for $\tau_{s}$ and $\psi_{s}$. Even closer at $10^{-10}$ away from the boundary, we find that the maximal disagreement in $\theta_{s}$, $u_{s}$, $\tau_{s}$, as well as $\psi_{s}$ is of order $10^{-3}$. This leads us to conclude that the error we make in the worst case is of the order of machine precision ($\approx 10^{-16}$) divided by the distance to the throat boundary times a factor of $10^{4}$.

### v2

@@ V.1.3 $\theta$-integral
With the solution for the polar angle, we proceed similarly to the previous subsection and state the final formula derived in [12] (the details can be found in Sec. III B there). The polar angle along the geodesic is given by
@@ V.1.4 $\psi$-integral
We can follow [12] and rewrite (52) as
@@
$\mathcal{I}_{\varphi,reg}=\nu_{r}\,\frac{a\,(\mathrm{ln}(r-r_{+})-\mathrm{ln}(r-r_{-}))}{r_{+}-r_{-}}\,.$ (56)
Furthermore, as shown in [12], (54) equates to
@@
and where $\gamma_{\pm}$ and $\mathcal{G}_{\theta}(\theta)$ are given by (46) and (47), respectively. However, we could not confirm the antiderivative of (38) presented in [12] when cross-checking it with numerical integration. Thus, we calculated $\mathcal{I}_{\pm}$ anew with the help of Mathematica [25]. The result can be found in (116)-(118) in Appendix E. The azimuthal angle of the source $\psi_{s}$ can be computed using (51) together with (56), (57) and (116). An example azimuthal trajectory can be found in Fig. 6.
@@ V.1.5 $u$-integral
As before, we can rewrite $I_{t}$ as in [12] to get
@@
$\mathcal{I}_{t,reg}(r)=$ $\nu_{r}\Big(r+\frac{a^{2}+r_{-}^{2}}{r_{-}-r_{+}}\,\mathrm{ln}(r-r_{-})$
@@
Next, we can use the solution of (63) as presented in [12], for which they found
@@
Lastly, we need the antiderivative of (39) for $\ell=0,1,2$. For $\ell=0$, the solution is (42). For $\ell=1,2$, respective formulas are presented in [12]. However, we could not confirm these results, and consequently, we calculated $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ anew. The results are presented in (117) and (118) in Appendix E. An example trajectory using (60) can be found in Fig. 7.
@@ V.2 Numerical solution
To solve this system numerically, we impose six initial conditions (two for each second-order equation and one for each first-order equation). The four initial values are the positional coordinates $(u_{o},\tilde{r}_{o},\theta_{o},\psi_{o})$ of the observer. We set the Mino time $\tau=0$ at the observer, thus we fix $\tilde{r}(0)=-\pi/2$, $\theta(0)=\theta_{o}$, $\psi(0)=\psi_{o}$, $u(0)=u_{o}$. For each second-order equation, we additionally need to state the rate of change of the coordinate at the observer, given by $\tilde{r}^{\prime}(0)=\nu_{r}$ and $\theta^{\prime}(0)=\beta$. We perform the numerical integration using Mathematica’s implementation of the explicit Bogacki-Shampine adaptive Runge-Kutta method [25]. Because we start the integration at the observer at $\tilde{r}_{o}=-\pi/2$, we will stop it once we reach the source at $\tilde{r}_{s}=\pi/2$. A comparison of the numerical and analytical solutions is presented in Appendix F.
@@ VI Visualizations
Using the presented solutions to the geodesic equations, we can inspect various vortical null geodesics in more detail. In the following, we consider a black hole with a rotation parameter $a/m=0.99$ and integrate from the position of the observer $(u_{o}=0,\,r_{o}=-\infty,\,\theta_{o}=\pi/4,\,\psi_{o}=0)$ to a test source located at $r_{s}=+\infty$. These values are chosen so that the inner throat remains significantly large while maximizing the asymmetry due to the rotation of the central object. Lowering the value $a/m$ only decreases the size of the observer’s field of view and renders the effect of the rotation on the null geodesics less prominent while not altering the overall qualitative behavior of null geodesics considered in this section. Note that by inverting the parameter $\tau$ along the geodesic, we obtain a trajectory that originates at the source and terminates at the observer.
In this section, for presentation, we use Kerr-Schild coordinates, which have an underlying flat metric, and we treat the coordinates $(x,y,z)$ as pseudo-Cartesian coordinates. Furthermore, going from one asymptotically flat region to the other (i.e., from $r>0$ to $r<0$ or vice versa) requires traversing the disk enclosed by the ring singularity [5]. In Kerr-Schild coordinates, the vortical null geodesics we consider can cross over from positive to negative values of $z$ only by traversing the disk in the $x$-$y$-plane, thereby changing the sign of $r$. Thus, in the following plots, the region with positive/negative $z$ corresponds to the region with positive/negative radius $r$. These two regions correspond to the northern hemisphere in their respective asymptotically flat ends of the spacetime (defined by the polar angle $\theta\in[0,\pi/2]$). Consequently, the horizons are only drawn at positive $z$ while the causality-violating region is only present at negative $z$.
@@ VI.1.2 Geodesics close to the inner throat boundary
Here, we address null geodesics close to the inner throat boundary (approximately $m\cdot 10^{-10}$ away). A geodesic close to the left vertex is shown in Fig. 9. Its main feature is that it remains close to the ring singularity in the region with negative $r$ for a significant portion of its journey. This period features several revolutions around the axis of symmetry and multiple polar oscillations between $\theta_{+}$ and $\theta_{-}$. However, it never crosses the causality-violating region.
For the right vertex of the inner throat, a corresponding trajectory in Kerr-Schild coordinates is presented in Fig. 10. The $\psi$-coordinate of this geodesic is not monotonically increasing or decreasing. In the exterior of a Kerr black hole, the order of null geodesics can be defined as the number of times it crosses the equatorial plane from the emitter to the observer [8]. Equivalently, one can consider the number of turning points in the polar angle along the trajectory [16]. Only the latter of these two definitions is meaningful for vortical null geodesics, which can never cross the equatorial plane. With this, the geodesic in Fig. 9 is of order four, while the one in Fig. 10 is of order three.
Furthermore, the geodesic close to the right vertex comes close to the ring singularity (up to $10^{-4}m$) but can never hit it [5]. However, it traverses the causality-violating region. To see this, we inspect the value of the metric component $g_{\psi\psi}$ along the geodesic. Fig. 11 shows the resulting plot. Since there are regions with $g_{\psi\psi}<0$, this geodesic passes through the causality-violating region multiple times along its path.
@@ VI.2.1 Field of view
With our solutions to the equations of motion for vortical null geodesics, we can furthermore answer the question of what portion of the sky at $r_{s}=+\infty$ is visible for an observer at $(u_{o}=0,\,r_{o}=-\infty,\,\theta_{o}=\pi/4,\,\psi_{o}=0)$. To this end, we discretize the inner throat, resulting in almost $1.5\cdot 10^{6}$ pairs of impact parameters $(\alpha,\,\beta)$ with the closest distance to the inner throat boundary of $3m\cdot 10^{-4}$. We calculate $\theta_{s}$ and $\psi_{s}$ for each of these parameters using our analytic solutions. To better track these pairs, we color-code them based on their distance to the center of the inner throat – warm colors are in the middle, while cool colors correspond to the region close to the inner throat boundary (see Fig. 12(a)).
@@
A further feature of the resulting curves in the polar plot is that the curves with a constant distance to the inner throat boundary develop caustic-like structures before they fold in on themselves at specific distances. Due to the overlap of curves, this is not visible in Figs. 12(b) and 12(c). However, it can be observed in a short animation available at [15]. We decrease the distance to the inner throat from frame to frame, thus approaching the boundary until the analytic and numerical solutions cease to coincide up to an acceptable margin (cf. Appendix F). It is visible that the field of view shifts in the anticlockwise direction in the polar plot as the corresponding curve in the inner throat approaches the boundary, thus covering large portions of the sky at $r_{s}=+\infty$. The fact that the curve in the polar plot folds in on itself results in the field of view covering the same part of the sky more than once. Consequently, the observer perceives multiple images of the same radiating object at $r_{s}=+\infty$ (as in the exterior case of a Kerr black hole [16]).
@@ VI.2.2 Observing the sky
A similar question concerns the appearance of a fixed test light source at $r_{s}=+\infty$ (e.g., the celestial sphere around the black hole) as seen by an observer at $(u_{o}=0,\,r_{o}=-\infty,\,\theta_{o}=\pi/4,\,\psi_{o}=0)$. For an initial study of the distortion effects (rotation, reflection, stretching, or compression of the image), we consider two sample light sources illuminating the sky at $r_{s}=+\infty$. We argue in Appendix G that we do not need to consider any gravitational redshift when the photons are emitted at $r_{s}=+\infty$ and observed at $r_{o}=$-$\infty$.
@@ VII Summary
An analysis of the angular potential showed that there are, at most, two distinct geodesics with a constant polar angle inside the inner throat. One is always visible and corresponds to the principal null direction on which the Eddington-Finkelstein-like coordinates are based. As such, it also has constant $\psi$ and $u$. The other one is only inside the inner throat for specific values of the rotation $a/m$ and the observer’s polar angle $\theta_{o}$. Moreover, vortical null geodesics can never cross the equatorial plane $\theta=\pi/2$. Consequently, an observer at $r_{o}=$-$\infty$ will only see at most half of the celestial sphere at $r_{s}=+\infty$. In addition, there is a further forbidden region around the equatorial plane from which such an observer receives no photons, see Fig. 3 and [23].
To visualize individual trajectories and simulate the observer’s view, we solved the geodesic equations of motion for vortical null geodesics inside the inner throat in Eddington-Finkelstein-like coordinates analytically and numerically, with the numerical integration providing an independent validation of the analytical solution. This way, we could correct some formulas in [12].
@@ Acknowledgements.
We thank Piotr Chruściel for valuable discussions and comments. This work is based on the Master’s thesis of the second author [23]. This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/P36455, 10.55776/Y963, 10.55776/COE1].
@@ Appendix A $\psi$ and $u$ for Principal Null Geodesics
As a geodesic with $(\alpha=+a\sin\theta_{o},\,\beta=0)$ corresponds to the principal null congruence on which the transformation to Eddington-Finkelstein-like coordinates is based on (cf. [22]), we claim that along this geodesic also $\psi$ and $u$ remain constant, which we prove below. Let us begin by evaluating the constants of motion (23) at these impact parameters, leading to
@@ Appendix B Definitions for the antiderivatives
To calculate the antiderivatives in Sec. V.1, we define additional quantities similar to [12].
In the case of no radial turning point, all roots of the radial potential $R(r)$ are complex with $r_{1}=r_{2}^{*}$ and $r_{3}=r_{4}^{*}$ (where the star denotes complex conjugation). These roots are given by [12]
@@
$h$ $=\mathrm{sign}(\cos\theta_{s})\,$, (100)
where $a_{1},b_{1},a_{2},b_{2}$ are uniquely specified by the restrictions $r_{1}=r_{2}^{\ast}$ and $r_{3}=r_{4}^{\ast}$ on the roots. For the calculation of the radial solution, we additionally need
@@ Appendix C Elliptic Integrals
For completeness, this section provides a basic introduction to elliptic functions following [1] and [10]. An integral of the form $\int R(x,y)\,\mathrm{d}x$ is called an elliptic integral if $R(x,y)$ is a rational function and $y=\sqrt{P(x)}$ where $P(x)$ is a polynomial either cubic or quartic in $x$. Generally, these integrals cannot be solved using elementary functions. All elliptic integrals can be reduced to integrals of elementary functions and three canonical elliptic integrals:
@@ Appendix E Antiderivatives of $I_{\pm}$, $I_{1}$, $I_{2}$
Here, we present the antiderivatives $I_{\pm}$, $I_{1}$, and $I_{2}$, defined in (38) and (39). We calculated these anew because we could not confirm the formulas presented in [12] when comparing them with numerical methods.
@@ Appendix F Comparing analytics and numerics
Far inside the inner throat (up to a distance of $m\cdot 10^{-3}$ from the inner throat boundary), the difference between the analytical and the numerical solutions for $\theta_{s}$ and $u_{s}$ is less than $10^{-9}$, while for $\tau_{s}$ and $\psi_{s}$ it is less than $10^{-8}$. We interpret this as a complete agreement of the two solutions. Fig. 14 shows the difference of the numerical and analytic solution along a vortical null geodesic with impact parameters $(\alpha=-0.15,\,\beta=0.1)$, which corresponds to the example presented in the figures in Sec. V.1. As expected, the error grows with decreasing distance to the throat boundary. At a distance of $m\cdot 10^{-8}$ from the boundary the maximal error for $\theta_{s}$ and $u_{s}$ is around $10^{-5}$, and $10^{-4}$ for $\tau_{s}$ and $\psi_{s}$. Even closer at $m\cdot 10^{-10}$ away from the boundary, we find that the maximal disagreement in $\theta_{s}$, $u_{s}$, $\tau_{s}$, as well as $\psi_{s}$ is of order $10^{-3}$. This leads us to conclude that the error we make in the worst case is of the order of machine precision ($\approx 10^{-16}$) divided by the distance to the throat boundary times a factor of $10^{4}$.